The vapour density of a mixture containing NO2 and N2O4 is 38.3 at 300K. The number of moles of NO2 in 100g of the mixture is  … ?     

Grade:

4 Answers

Sunil Kumar FP
9 years ago
The vapor density of a gas is the density of that gas relative to that of hydrogen gas.

Therefore,

Vapor density = (density of NO2/N2O4 mixture)/ (density of H2 gas)
Vapor density = (mass of NO2/N2O4 mixture)/ (mass of H2 gas)
Vapor density = (molar mass of NO2/N2O4 mixture)/ (molar mass of H2 gas)
Vapor density = (molar mass of NO2/N2O4 mixture)/ 2
Molar mass of NO2/N2O4 mixture = Vapor density * 2
= 76.6 g mol^(-1)

Let the percentage of NO2 gas be X, the percentage of N2O4 gas be 1-X

X(molar mass of NO2 gas) + (1-X)(molar mass of N2O4) = 76.6
X(46) + (1-X)(92) = 76.6
X = 33.48%

Mass of NO2 in 100g of mixture = 33.48g
Number of moles of NO2 gas = 33.48/46 = 0.728 mol

Nandariya Ajit
7 years ago
Mixture molecule waite=2×38.3=76.6Number of mixture mole=100/76.6100 gm mixture in x gm NO2NO2 number of mole=x/46 N2O4 number of mole=100-x/92x/46+(100-x)/92=100/76.6 =20.1 gmNumber of NO2 mole =20.1/46=0.437 moleAns=0.437
Kriti Sharma
5 years ago
vapour density= 38.3
mass of 1 mole of mixture= vapour density * 2= 3.38*2=76.6
no of moles in the mixture= 100/76.6=1.305 moles
now, 
      moles of NO2 + moles of N2O4= 1.305
      let the mass of NO2 be x
      x/46 + 100-x/92 = 1.305
      2x+100-x/92 = 1.305
      x+100 = 120.06
      x= 20.1
so, 20.1 grams of NO2 is present in 100 g of solution 
no of moles of NO2= 20.1/46=4.37 moles
 
Ajeet Tiwari
3 years ago
hello students

solution is as follows

Mol mass of mixture=2×38.3=76.6

No. of moles in 100 g of mixture=76.6100​

Let a g ofNO2​is present in mixture.

Molesof NO2​+Molesof N2​O4​=Molesofmixture

46a​+92100−a​=76.6100​"or"a=20.10g

Moles ofNO2​in mixture=20.10/46 ​=0.437


Hope it helps
Thankyou and Regards